OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1

OpenStudy (anonymous):

D. an = -4 • an-1

OpenStudy (anonymous):

\[r = \frac{a_n}{a_{n-1}} = \frac{a_2}{a_1} = -4\]the recursive formula is\[a_n = ra_{n-1}\]\[a_n = -4a_{n-1}\]